# Quantum complexity of topological phases. Or lack thereof

### v2

@@ Abstract
Topological phases are expected to carry enhanced computational power, stemming from the long-range entanglement that characterizes them. Here we investigate whether this advantage is reflected in quantum magic, quantified through stabilizer Rényi entropies (SREs). We analyze one-dimensional models with an exact duality between a symmetry-protected topological phase and a trivial one. Surprisingly, a finite (parameters dependent) asymmetry appears only when boundary conditions break the duality, and emerges as well in the transverse-field Ising chain, which lacks topological order. These results complement recent analysis in the literature and show that the total magic of a state is unable to detect topological phases, indicating either that a different property is required to capture their unique quantum contribution, or that not all forms of long-range entanglement constitute a genuine computational resource.
One of the central challenges in quantum information science is to identify and characterize the fundamental properties that a quantum state must possess in order to enable operations that outperform classical computation — in other words, to achieve quantum computational advantage or quantum supremacy [1]. For a long time, the characterization of these states focused predominantly on their entanglement properties [2, 3]. However, over time, it became increasingly evident that entanglement alone does not provide a complete picture. Notably, certain highly entangled states can still be efficiently simulated on classical computers [4], while some computational models exhibit a quantum advantage even in the absence of significant entanglement [5]. This realization has led to the identification and study of several additional resources over the years [6]. Among them, the so-called quantum nonstabilizerness, also known as quantum magic [7, 8, 9, 10], has acquired a relevant role in the past years. Quantum magic quantifies the extent to which a state departs from the stabilizer manifold, i.e., the manifold of the states that can be obtained from a reference one using circuits containing only Clifford gates [11], a subset of quantum operations that can be efficiently simulated on a classical computer. In this sense, magic captures a form of non-classicality necessary for universal quantum computation.
This issue is intimately connected to the study of quantum many-body systems, whose ground states often serve as natural candidates for implementing a wide range of quantum information processing tasks. As a result, quantum magic is emerging as an operational criterion to distinguish phases of matter that are classically simulable from those that support a genuine quantum advantage [12, 13, 14], much like entanglement has already done [15, 16, 17, 18, 19]. Among others, topological phases of matter are widely expected to possess significant computational power due to their robustness against local perturbations and their ability to encode quantum information non-locally [20]. These phases exhibit exotic correlations not captured by local order parameters and can support computational models such as topological quantum computation, where logical gates are implemented through the braiding of anyonic excitations [21, 22]. In one-dimensional systems, the topological order is more fragile, but a weaker analog exists in the form of symmetry-protected topological phases (SPTPs) [23, 24]. Examples include the one-dimensional cluster-Ising model [25, 26, 27, 28] and the dimerized XX chain [29, 30], which can be obtained from the Su–Schrieffer–Heeger (SSH) model [31, 32, 33] via a Jordan–Wigner transformation. These systems display non-trivial entanglement structures and symmetry-protected edge modes; as long as symmetry-breaking perturbations are excluded, they retain their topological character and can outperform trivial phases. In particular, they support non-trivial forms of quantum computation, particularly in measurement-based schemes [34, 35]. Most importantly, as argued in [36, 23] and elaborated in [37, 38], the long-range entanglement of the SPTPs is a precious resource that cannot be generated from trivial ones without implementing circuits whose depth diverges with the system’s size. Thus, only non-trivial topological phases can serve as universal quantum simulators. If computational complexity classes are equivalent, this implies that SPTPs should carry a higher intrinsic computational complexity, measurable as an enhanced amount of quantum magic [8, 7]. It should be remarked, however, that recent works have realized that by allowing non-unitary processes, such as measurements, in a circuit, local entanglement can be promoted to global entanglement [39, 40]. In this way, (abelian) topological order can be generated using much shallower circuits, thus providing a potential loophole to the presumed primacy of topological order [41, 42, 43, 44, 45].
In this letter, we test the relation between complexity and topological order by calculating the quantum magic of one-dimensional SPTPs and comparing it with that of their topologically trivial counterparts. We focus on models that possess a duality transformation connecting a SPTP to a trivial one, and find that their magic coincides, except when boundary conditions force a violation of the duality symmetry. In such cases, we observe a finite, parameter-dependent difference in quantum magic that originates from boundary effects that tend to be independent of the system size in the thermodynamic limit. The inability of the total magic to detect the long range order of a state highlights that a single number is inefficient in characterizing the various resources hidden in a state complexity. Indeed, recent studies have also focused on long-range nonstabilizerness (LRM), that is, the magic that cannot be removed by local, shallow circuits [46, 47, 48] and found that topological order comes in different fashions, which may or may not possess LRM. However, it should be pointed out that such classifications arise within a circuit-based approach, which connects the circuit simulability with an intrinsic property of a state. Such a computational point of view is indeed of growing interest in the community, but in this work we are more interested in directly assessing a state’s complexity and what it can tell us.
@@
Figure 2: Behavior of $\Delta\mathcal{M}_{2}(\delta)=\mathcal{M}_{2}(\delta)-\mathcal{M}_{2}(-\delta)$, as function of $\delta>0$ for the model in Eq. (2). For small $L$, $\Delta\mathcal{M}_{2}$ shows a size dependence, but as $L$ increases the curves start to converge to a non-constant limiting function of $\delta$. This finite, parameter-dependent asymmetry is inconsistent with the behavior of a genuine topological invariant. The self-dual point $\delta=0$ is excluded, since $\Delta\mathcal{M}_{2}=0$ by definition.
To isolate the potential topological contribution to SRE, for each $L$ we define the difference in magic between the ground states at symmetric points with respect to the critical point $\Delta\mathcal{M}_{2}(\delta)=\mathcal{M}_{2}(\delta)-\mathcal{M}_{2}(-\delta).$ These differences are shown in Fig. 2. For $\delta>0$ $\Delta\mathcal{M}_{2}(\delta)$ remains consistently positive and rapidly converges to a finite value as the system size increases. For $L\geq 30$, the curves start to collapse onto a single, well-defined limiting function except close to $\delta=0$, as the correlation length becomes comparable with the system sizes. However, this limiting function retains a non-trivial dependence on $\delta$, suggesting that the asymmetry in quantum magic does not have a topological origin. In fact, topological quantities are expected to be quantized and constant within a given macroscopic phase [58, 59, 60, 61, 13], whereas our results feature a smooth dependence on the microscopical parameters.
@@
Furthermore, consistent with the findings for the dimerized XX model, when the boundary conditions are changed from open to periodic – closing the loop by adding to the Hamiltonian in Eq. (3) the terms $\sigma_{L-1}^{x}\sigma_{L}^{z}\sigma_{1}^{x}-\sigma_{L}^{x}\sigma_{1}^{z}\sigma_{2}^{x}+\lambda\sigma_{1}^{y}\sigma_{L}^{y}$ – the quantity $\Delta\mathcal{M}_{2}(\lambda)$ vanishes identically. PBCs restore the exact duality $\lambda\leftrightarrow 1/\lambda$ and remove the possibility of observing edge states, but they do not destroy the topological nature of the $\lambda<1$ phase, which remains signaled by a finite string-order parameter [26]. Thus, exactly dual points have the same SRE (since, once again, the duality trasformation is of the Clifford type), demonstrating the inability of quantum magic to detect the presence of topological order. Once again, the difference in SRE arises solely from the violation of duality symmetry at the boundaries. We remark that this is in line with the results of [48], where it was proven that in 1D cluster states have high computational complexity from the point of view of a Montecarlo simulation (due to their sign problem structure), but not from a Clifford resource point of view.
@@
As can be appreciated from Fig. 5, the behavior of quantum magic $\mathcal{M}_{2}$ for the ground state of the transverse field Ising model is qualitatively the same as that of the previous models: with OBCs, $\mathcal{M}_{2}$ has a peak inside the ferromagnetic phase, which grows and moves towards $h=1$ as the system size is increased. This behavior is sensibly different from what was obtained in [12], where the same model was analyzed assuming PBCs, i.e. with exact dual symmetry, and agrees with what we get under the same conditions: the maximum is always centered at $h=1$. Moreover, the quantity $\Delta\mathcal{M}_{2}(h)=\mathcal{M}_{2}(h)-\mathcal{M}_{2}(1/h)$ ($h\in[0,1)$) vanishes identically when the duality symmetry holds, independently of $h$ and $L$ because the Kramers-Wannier duality is magic-preserving. Instead, as also observed for the other models, breaking the duality induces an asymmetry in the SRE. As shown in Fig. 6, with OBCs $\Delta\mathcal{M}_{2}(h)$ vanishes only in the limit of $h\rightarrow 0$. Otherwise, its value depends on $L$ but the different curves tend to collapse as soon as the dimension of the system becomes greater than the correlation length. Obviously, since no topological phase is present in this model, the existence of such a difference cannot be justified by the presence of topological order.
@@
Before concluding, we emphasize that all evaluations of quantum magic in this Letter were carried out in the computational basis. Since magic is inherently basis-dependent, one may wonder whether our results are tied to this choice or hold more generally. A possible answer to this question is that attempted for instance in [68, 46, 47, 48]: minimizing magic through local unitaries in a circuit-based approach. Indeed in this way one aims at extracting the “most interesting” part of non-stabilizerness. As we are instead looking at the whole complexity of a state, we just tested local unitaries composed of independent single-spin rotations. For small systems ($L\leq 10$), we found that our framework remains robust: with PBCs, the magic of dual points is identical in any basis, while opening the chain, breaking the duality and inducing the formation of edge states breaks this symmetry. Hence, the physical conclusions presented here are not restricted to a specific basis but acquire a general validity. Finally, from our results we recover that, the computational basis minimizes the magic for all models and parameter choices, and thus coincides with the non-local $L$-partite magic [69, 70, 71].
In conclusion, we calculated and analyzed the rank $2$ Stabilizer Renyi Entropy $\mathcal{M}_{2}$ to look for the contribution that topological order adds to quantum magic. To isolate this element, we focused on models that have a duality symmetry between a SPTP and a trivial one, such as the dimerized XX model, which is directly mappable onto the SSH model, and the Cluster-Ising chain. The long-range entanglement that characterizes SPTPs is known to be a precious resource that is expected to provide a quantum advantage [36, 23], and accordingly we expected an increase in SRE between dual points. Surprisingly, however, we found that they exhibit exactly the same amount of magic. A sub-extensive asymmetry appears only under open boundary conditions, which break the exact duality and introduce small corrections to the SRE. This finite difference is not a constant throughout the phase and as a final proof of its non-topological origin we checked that the quantum Ising chain presents the same phenomenology. These results are summarized in Table 1.
Our findings are rather unexpected and should be contrasted with the case of topological frustration, where a delocalized topological excitation is injected in the ground state by an appropriate choice of boundary conditions [72, 73, 61, 74]. In that case, the excitation contributes an additional amount of quantum magic that grows logarithmically with system size compared to the non-frustrated counterpart [13, 14]. The long-range entanglement generated by the topological excitation provides an added complexity, while apparently that of a conventional SPTP does not. Indeed, $\mathcal{M}_{2}$ measures the amount of non-Clifford resources needed to generate a given state [11]. Our results seem to imply that is possible to efficiently generate topological order using the same amount of resources needed for much simpler orders. Such a conclusion seemingly contradicts the expected complexity equivalence between different quantum computation schemes and needs to be understood better. Indeed, to the best of our knowledge, there is no resource theory that takes specific advantage of topological order, leveraging over adiabatic continuation as a free operation and measuring the contribution that crossing a phase transition or changing topology brings. To the contrary, recent results showing how Local Operations and Classical Communications (LOCC) can be utilized to convert local into global entanglement [39, 40, 41, 42, 43, 44, 45]., indicate that in non-unitary computational schemes, topological order is not a necessary resource. Nonetheless, the classification of these latter proposals within a Clifford resource theory is not straightforward, and thus it is not clear if our findings stem from such approaches or are indicative of something else. It should also be clarified if our results extend to other topological models, even in higher dimensions: in the absence of exact dualities, one could check whether the growth of $\mathcal{M}_{2}$ with the system size deviates from the one expected considering purely local order [49].
@@ Acknowledgements.
The research leading to these results has received funding from the following organizations: European Union via ICSC - Italian Research Center on HPC, Big Data and Quantum Computing (NextGenerationEU Project No. CN00000013), project EuRyQa (Horizon 2020), project PASQuanS2 (Quantum Technology Flagship); Italian Ministry of University and Research (MUR) via: Quantum Frontiers (the Departments of Excellence 2023-2027); the World Class Research Infrastructure - Quantum Computing and Simulation Center (QCSC) of Padova University; Istituto Nazionale di Fisica Nucleare (INFN): iniziativa specifica IS-QUANTUM. S. P. acknowledges financial support from the Ministero dell’Universit‘a e della Ricerca (MUR) and the Project PRIN 2022 number 2022W9W423 funded by the European Union Next Generation EU. SBK acknowledges support from the Croatian Science Foundation (HrZZ) Projects DOK-2020-01-9938.

### v3

# Resource complexity of Symmetry Protected Topological phases

@@ Abstract
We pursue the identification of quantum resources carried by topological order, by evaluating quantum magic, quantified through the rank-$2$ Stabilizer Rényi entropy $\mathcal{M}_{2}$, in one-dimensional systems hosting symmetry-protected topological phases (SPTP). Focusing on models with an exact duality between an SPTP and a trivial one, namely the dimerized XX and the Cluster–Ising chains, we show that dual points exhibit identical amounts of magic, even though they belong to distinct topological sectors. A subextensive asymmetry arises only under open boundary conditions, where edge effects break the duality, but this correction is non-topological and depends on microscopic parameters. These results stand in contrast to the case of topological frustration, where delocalized excitations enhance the magic logarithmically with system size. They also complement recent analyses in the literature, showing that the total magic is largely insensitive to the presence of topological order, hence suggesting that topological order is not necessarily a genuine computational resource.
preprint: RBI-ThPhys-2025-38
A key challenge in quantum information science is to identify the resources that a quantum state must possess to perform tasks beyond the reach of classical computation, that is, to establish quantum supremacy [1]. This question naturally connects to the study of quantum many-body systems, whose ground states often embody such resources. For a long time, entanglement [2, 3] was believed to be the main bridge between many-body physics and quantum information [4, 5, 6, 7, 8]. However, it has become clear that entanglement alone cannot fully account for quantum advantage. Some highly entangled states can be efficiently simulated using classical resources [9], while certain computational models exhibit quantum advantage even with little entanglement [10]. This insight has motivated a broader search for other operationally meaningful quantum resources [11]. Among these, quantum magic (or nonstabilizerness) [12, 13, 14, 15] has recently emerged as a key quantity, measuring how far a quantum state lies outside the stabilizer manifold. The latter is the set of states obtainable from computational-basis states through Clifford circuits [16, 17, 18], which themselves can be efficiently simulated on classical computers. Therefore, in loose terms, magic can be interpreted as the amount of non-Clifford resources needed to prepare a given quantum state starting from a computational basis state. It is thus unsurprising that, similarly to entanglement in the past, magic has become a powerful proxy linking phases of matter to their computational capabilities [19, 20, 21].
Among quantum phases, topological ones are especially relevant, as they provide intrinsic protection against local perturbations and encode information nonlocally [22]. They exhibit correlations beyond local order parameters, host long-range entanglement, and enable fault-tolerant computation through the braiding of anyonic excitations [23, 24]. Given these properties, one might expect topologically ordered phases to exhibit a higher complexity and thus host a large quantum magic, reflecting their highly nonclassical structure. However, this connection is not universal: the paradigmatic toric code [23], though topologically ordered, admits as ground state a stabilizer state with no magic [17, 18].
In one-dimensional systems, the topological order is more fragile. Still, a weaker analog exists in the form of symmetry-protected topological phases (SPTPs) [25, 26], that are phases that present topological properties only as long as symmetry-breaking terms are excluded. Examples include the one-dimensional cluster-Ising model [27, 28, 29, 30] and the dimerized XX chain [31, 32], which can be obtained from the Su–Schrieffer–Heeger (SSH) model [33, 34, 35] via a Jordan–Wigner transformation. Compared to inherent topological phases, topological entanglement entropy [36, 37] is not uniquely defined in SPTPs. For instance, under open boundary conditions, the emergence of edge states gives rise to nonvanishing edge-to-edge entanglement [38, 39]. Similarly, only the weaker form of symmetry-preserved long-range entanglement applies, and a finite-depth unitary circuit can remove all entanglement if the symmetry is not enforced, whereas this is not possible for true topological phases [40, 25, 41]. The distinction between these two categories becomes less sharp when shallow circuits are supplemented by measurement-based protocols, as it appears that certain types of inherent topological order phases can be created with finite-depth circuits [42, 43]. Nevertheless, SPT phases present relevant properties, even from a computational point of view. In particular, one-dimensional SPT ground states (e.g., the cluster state) equip quantum wires for measurement-based quantum computation (MBQC) [44, 45, 46], while universal MBQC requires higher-dimensional resource states [47, 48, 49, 50, 51, 52, 53, 54].
In this Letter, we explore the connection between computational complexity and topological phases by evaluating the quantum magic of one-dimensional SPTPs and comparing it with that of their topologically trivial counterparts. To make this comparison stringent, we focus on models possessing an exact duality between an SPT and a trivial phase, namely the aforementioned dimerized XX and Cluster–Ising chains. We find that dual points exhibit identical amounts of magic whenever the duality is preserved (e.g., under periodic boundary conditions), whereas a subextensive asymmetry arises only with open boundaries, where edge effects explicitly break duality. However, this correction is not of topological origin, as it depends on microscopic Hamiltonian parameters, and, as we will show in the following, a similar phenomenology also occurs in the quantum Ising chain. Our findings, therefore, suggest that assessing computational complexity may require considering a different quantity, such as the long-range magic (LRM) [46, 55, 56], which captures the component of nonstabilizerness that cannot be eliminated by shallow local circuits.
@@
Figure 2: Behavior of $\Delta\mathcal{M}_{2}(\delta)=\mathcal{M}_{2}(\delta)-\mathcal{M}_{2}(-\delta)$, as function of $\delta>0$ for the model in Eq. (2). For small $L$, $\Delta\mathcal{M}_{2}$ shows a size dependence, but as $L$ increases, the curves start to converge to a non-constant limiting function of $\delta$. This finite, parameter-dependent asymmetry is inconsistent with the behavior of a genuine topological invariant. The self-dual point $\delta=0$ is excluded, since $\Delta\mathcal{M}_{2}=0$ by definition.
To isolate the potential topological contribution to SRE, for each $L$, we define the difference in magic between the ground states at symmetric points with respect to the critical point $\Delta\mathcal{M}_{2}(\delta)=\mathcal{M}_{2}(\delta)-\mathcal{M}_{2}(-\delta).$ These differences are shown in Fig. 2. For $\delta>0$ $\Delta\mathcal{M}_{2}(\delta)$ remains consistently positive and rapidly converges to a finite value as the system size increases. For $L\geq 30$, the curves start to collapse onto a single well-defined limiting function except close to $\delta=0$, as the correlation length becomes comparable with the size of the system. However, this limiting function retains a non-trivial dependence on $\delta$, suggesting that the asymmetry in quantum magic does not have a topological origin. In fact, topological quantities are expected to be quantized and constant within a given macroscopic phase [39, 66, 67, 68, 20], whereas our results feature a smooth dependence on the microscopical parameters.
@@
Furthermore, consistent with the findings for the dimerized XX model, when the boundary conditions are changed from open to periodic – closing the loop by adding to the Hamiltonian in Eq. (3) the terms $\sigma_{L-1}^{x}\sigma_{L}^{z}\sigma_{1}^{x}-\sigma_{L}^{x}\sigma_{1}^{z}\sigma_{2}^{x}+\lambda\sigma_{1}^{y}\sigma_{L}^{y}$ – the quantity $\Delta\mathcal{M}_{2}(\lambda)$ vanishes identically. PBCs restore the exact duality $\lambda\leftrightarrow 1/\lambda$ and remove the possibility of observing edge states, but they do not destroy the topological nature of the $\lambda<1$ phase, which remains signaled by a finite string-order parameter [28]. Thus, exactly dual points have the same SRE (since, once again, the duality transformation is of the Clifford type), demonstrating the inability of quantum magic to detect the presence of topological order. Once again, the difference in SRE arises solely from the violation of duality symmetry at the boundaries. We remark that this is in line with the results of [56], where it was proven that in 1D cluster states have high computational complexity from the point of view of a Montecarlo simulation (due to their sign problem structure), but not from a Clifford resource point of view.
@@
As can be appreciated from Fig. 5, the behavior of quantum magic $\mathcal{M}_{2}$ for the ground state of the transverse field Ising model is qualitatively the same as that of the previous models: with OBCs, $\mathcal{M}_{2}$ has a peak inside the ferromagnetic phase, which grows and moves towards $h=1$ as the system size is increased. This behavior is sensibly different from what was obtained in [19], where the same model was analyzed assuming PBCs, i.e., with exact dual symmetry, and agrees with what we get under the same conditions: the maximum is always centered at $h=1$. Moreover, the quantity $\Delta\mathcal{M}_{2}(h)=\mathcal{M}_{2}(h)-\mathcal{M}_{2}(1/h)$ ($h\in[0,1)$) vanishes identically when the duality symmetry holds, independently of $h$ and $L$ because the Kramers-Wannier duality is magic-preserving. Instead, as also observed for the other models, breaking the duality induces an asymmetry in the SRE. As shown in Fig. 6, with OBCs $\Delta\mathcal{M}_{2}(h)$ vanishes only in the limit of $h\rightarrow 0$. Otherwise, its value depends on $L$, but the different curves tend to collapse as soon as the dimension of the system becomes greater than the correlation length. Obviously, since no topological phase is present in this model, the existence of such a difference cannot be justified by the presence of topological order.
@@
Before concluding, we emphasize that all evaluations of quantum magic in this Letter were carried out in the computational basis. Since magic is inherently basis-dependent, one may wonder whether our results are tied to this choice or hold more generally. Given the difficulty of evaluating magic, a comprehensive analysis over all possible bases in all possible systems is not feasible. Nevertheless, by restricting our study to small systems ($L\leq 10$) and to cases where the change of basis is implemented through unitaries composed of independent single-spin rotations, such an analysis becomes tractable. Our framework demonstrates notable robustness: under periodic boundary conditions (PBCs), the magic at the dual points remains identical across all bases, whereas opening the chain—thus breaking duality and inducing the emergence of edge states—breaks this symmetry. Hence, the physical conclusions presented here are not restricted to a specific basis but acquire a general validity. Finally, from our results, we recover that the computational basis minimizes the magic for all models and parameter choices, and thus coincides with the non-local $L$-partite magic [75, 76, 77].
In conclusion, we have analyzed the rank-$2$ Stabilizer Rényi Entropy $\mathcal{M}_{2}$ to investigate the possible contribution of topological order to quantum magic in 1D SPTPs. To isolate it, we focused on models exhibiting an exact duality, namely the dimerized XX model and the Cluster–Ising model. Since SPTPs are generally regarded as valuable quantum resources expected to yield computational advantage within the MBQC framework, an enhancement of $\mathcal{M}_{2}$ between dual points might be anticipated. On the contrary, we find that the dual phases exhibit identical amounts of magic. A subextensive asymmetry appears only under open boundary conditions, where the duality is broken by edge effects, introducing finite corrections to $\mathcal{M}_{2}$. This difference is not constant across the phase and, as confirmed by the identical behavior observed in the quantum Ising chain, is of non-topological origin. These results are summarized in Table 1. Our findings stand in stark contrast to the case of topological frustration in one dimension [78, 79, 68, 80], where the introduction of a delocalized topological excitation into the ground state leads to an additional contribution to quantum magic that grows logarithmically with system size [20, 21].
Since $\mathcal{M}_{2}$ quantifies the amount of non-Clifford resources required to generate a given state, our results indicate that the ground states of SPTPs can be efficiently prepared using the same resources as their topologically trivial counterparts. However, given that these phases differ fundamentally within the MBQC paradigm, this observation appears to challenge the expected equivalence of computational complexity across different quantum computation models and thus calls for a deeper understanding. On the other hand, our findings are consistent with recent results showing that Local Operations and Classical Communication (LOCC) can convert local into global entanglement [42, 43, 81, 82, 83, 84, 85], suggesting that even intrinsic topological order may not be a necessary resource in non-unitary computational schemes. Nevertheless, the classification of these schemes within a Clifford resource framework remains nontrivial, and it is therefore unclear whether our results are a manifestation of such mechanisms or evidence of a distinct underlying phenomenon.
To clarify the generality of our findings they should be extended to truly topological models, with both abelian and non-abelian symmetries, in higher dimensions: in absence of exact dualities, one could check whether the growth of $\mathcal{M}_{2}$ with the system size deviates from the one expected considering purely local order [57] to assess whether topological order carries or not an intrinsic amount of quantum complexity.
@@ Acknowledgements.
The research leading to these results has received funding from the following organizations: European Union via ICSC - Italian Research Center on HPC, Big Data and Quantum Computing (NextGenerationEU Project No. CN00000013), project EuRyQa (Horizon 2020), project PASQuanS2 (Quantum Technology Flagship); Italian Ministry of University and Research (MUR) via: Quantum Frontiers (the Departments of Excellence 2023-2027); the World Class Research Infrastructure - Quantum Computing and Simulation Center (QCSC) of Padova University; Istituto Nazionale di Fisica Nucleare (INFN): iniziativa specifica IS-QUANTUM. A.G.C. acknowledges support from the MOQS ITN programme, a European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement number 955479. G. T., F. F., and S. M. G. acknowledge support of “Implementation of cutting-edge research and its application as part of the Scientific Center of Excellence for Quantum and Complex Systems, and Representations of Lie Algebras”, Grant No. PK.1.1.10.0004, co-financed by the European Union through the European Regional Development Fund - Competitiveness and Cohesion Programme 2021-2027. S. B. K. acknowledges support from the Croatian Science Foundation (HrZZ) Projects DOK-2020-01-9938. S. P. acknowledges financial support from the Ministero dell’Universitá e della Ricerca (MUR) and the Project PRIN 2022 number 2022W9W423 funded by the European Union Next Generation EU.